OpenStudy (anonymous):

the weight of W of an object on the moon varies directly as the weight E on earth a person who weighs 163lbs on earth weighs 32.6lb on the moon how much would a 142lb person weigh on the moon?

OpenStudy (anonymous):

set up a proportion 163/32.6=142/x solve for x.

OpenStudy (amistre64):

Mw = k * Ew 32.6 = k 163 k = 32.6/163 = 2/10 = 1/5 Mw = 1/5 Ew Mw = 1/5 142 Mw = 28.4 lbs

OpenStudy (anonymous):

thanks i stepped away for awhile